OpenStudy (anonymous):

find tan4x? tanx=1/5

OpenStudy (anonymous):

what's the question mark for?

OpenStudy (anonymous):

shfit tan=1/5 ,then u will know x = ? ,then sub. the x in tan4x :D

OpenStudy (asnaseer):

use the formula:\[\tan(2x)=\frac{2\tan(x)}{1-\tan^2(x)}\]to first find tan(2x) and then tan(4x)

OpenStudy (asnaseer):

Once you know tan(2x), use same formulae to get:\[\tan(4x)=\frac{2\tan(2x)}{1-\tan^2(2x)}\]